Solve this system of equations:
y=x^2-2x-1
y=-x^2+4x-1

Answer by farbonara(13)

y=x^2-2x-1 ---(1)
y=-x^2+4x-1 ---(2)
(1)=(2)
x^2-2x-1=-x^2+4x-1
2x^2-6x=0
x^2-3x=0
(x)(x-3)=0
x=0 or x=3
Using (1),
When x=0,y=-1
When x=3,y=2